# Re-Simulation-based Self-Supervised Learning for Pre-Training Foundation Models

### v1

@@ 2.3.1 In-distribution classification and robustness
Figure 5: Tagging performance on the Higgs vs. QCD classification task. All numbers are obtained on a sample comprising only jets from the default parton shower scenario. Fine-tunings on top of a RS3L space that was trained with 5M jets are shown in light and dark blue, for graphs that have floating and fixed weights, respectively. The number in parentheses indicates the fine-tuning dataset size. Fully-supervised networks with various training sizes are shown in shades of grey. Three repeated trainings with different random initialization are used to compute an average and standard deviation for each ROC. These are shown as bands. Note that fixed-weight trainings do not have any appreciable uncertainty.
@@
The most straightforward application of RS3L is a binary classification fine-tuning to separate the same jet classes that are used to train the RS3L space. Thus, we consider the canonical HEP classification task of discriminating between jets originating from Higgs bosons and jets originating from QCD, a prerequisite for the precise understanding of Higgs boson couplings to bottom quarks [38, 39, 40, 41, 42].
@@
Two variants are considered for fine-tuning the RS3L space: (i) The weights of the backbone model are kept fixed, and a task-specific head consisting of a one-layer multilayer perceptron (MLP) is added, whose parameters are trained on the task at hand. (ii) The weights of the backbone model are used as initial parameters but are allowed to be adjusted (``floated'') together with the ones of the task head.
@@
A comparison of RS3L spaces fine-tuned for a Higgs vs. QCD classification task is presented in Figure 5, which shows the receiver operating characteristic (ROC) curve calculated from nominal jets only. The QCD rejection rate (1/ false positive rate) is shown in Table 1 for various Higgs tagging efficiencies, again computed from only nominal jets. Compared to a fully-supervised network trained on 8M jets, a fine-tuning using 3M jets on top of a 5M RS3L space yields a similar performance. However, if we instead match the yields of the training dataset of the fully-supervised approach and the fine-tuning (i.e., both training on 3M labeled samples), a large gain in performance is observed for the RS3L + fine-tuning strategy. This indicates that a common, large pre-training allows for efficient fine-tuning on comparably much smaller dataset sizes than a fully supervised setup would use.
@@
Furthermore, it is important to note that all training configurations that use floating networks are more performant than networks trained with fixed weights on top of the RS3L backbone. This is to be expected because in the fixed weight case the model is severely restricted as the only trainable parameters belong to the MLP head. Lastly, all networks outperform $N_{2}$ by a wide margin.
@@
In addition to background rejection, it is necessary to examine the classifier under other metrics. In particular, we consider how the network response varies under different parton shower configurations used in the simulator (which we call the network's ``robustness''). The robustness is quantified by calculating the Wasserstein distance [43, 44, 45] between the one-dimensional distributions of the classifier output of nominal and, respectively, augmented jets. The distances are presented in Table 2. Each column shows the distance between the distribution for augmented jets (indicated by column title) and nominal jets. A lower Wasserstein distance indicates a network that is less sensitive to the showering configurations. As expected, the distance between nominal jets and jets re-showered with just a different numerical seed is smaller than the distance between nominal jets and any other augmentation (``FSR up'', ``FSR down'', or ``Herwig''). Especially the distance between nominal jets and Herwig7-showered jets is significant.
@@
Comparing the different training strategies, the most robust networks are trained using the RS3L method. With a fixed-weight approach, the Wasserstein distance is significantly smaller than for all other networks. This is a direct benefit of the RS3L pre-training strategy of minimizing the distance between positive pairs. Including Herwig7 augmentations in RS3L or, respectively, full supervision greatly reduces the distances compared to setups without Herwig7 augmentations. Interestingly, we observe that already including seed and/or FSR augmentations in RS3L helps fine-tuned networks to be more robust against Herwig7-induced domain shift compared to fully supervised strategies. Whether this indicates that a gain in robustness is potentially attainable simply by training the backbone with re-seeded jets merits further investigation that we leave to future studies.
@@
Table 2: Wasserstein distances between distributions of tagger output for nominal and augmented jets. Each column indicates the distances between the nominal and the respective augmentation. The number of training events is included for each training setup. Fine-tunings are performed on a 5M RS3L pre-training with either fixed or floating backbone weights. The inclusion of specific augmentations (e.g. ``seed'' or ``FSR'') in the training setup indicates that the pre-training and fine-tuning or supervision is performed on that set of augmentations. Uncertainties are derived from three independent trainings.
@@
2.3.2 Out-of-distribution classification task
We now shift to out- task to address the degree to which a general representation of showering jets was achieved through the RS3L pre-training. To investigate this, we fine-tune RS3L for a classification task that falls outside of the distributions used for pre-training, namely the task of discriminating between QCD jets and jets from ``hadronic'' W boson decays. Contrary to the Higgs boson decays, jets from W boson decays do not feature $B$ hadrons [^6] The authors note that the absolute tagging performance on the W vs. QCD and Higgs vs. QCD datasets cannot be compared directly because of the composition of the sample, see A. The experimental signature therefore falls out of the distribution of the Higgs jets used to train the RS3L backbone.
@@ 3 Outlook
Enforcing domain completeness through a strategy to embed the known modeling uncertainties within the self-supervised space directly maps this work to the quality of the simulation used at training. Improved simulators can lead to improved embeddings, which, in turn, lead to a better understanding of the underlying physics. As a result, R3SL can adapt to the next generation of simulation modeling that stands to be substantially better than the current simulation. Through self-supervised approaches, we can continually improve the quality of the physics that is extracted, provided we continue to improve the physics model in the samples used for the self-supervision.
In future work, alternative self-supervised learning approaches to the employed SimCLR strategy, such as VICReg [46], SimSiam [47], and BarlowTwins [48] should be explored to study the practical impact of precise components of their loss functions. Additionally, the dataset size for the pre-training should be varied to systematically study the regimes in which the performances of SSL, fine-tuning, and fully-supervised strategies may saturate.
@@ 4 Acknowledgments
BM acknowledges the support of the Alexander von Humboldt foundation and of Schmidt Futures. MK is supported by the US Department of Energy (DOE) under grant DE-AC02-76SF00515. PH and JK are supported by the Institute for Artificial Intelligence and Fundamental Interactions (IAIFI) under the NSF grant #PHY-2019786 and the Accelerated AI Algorithms for Data Driven Discovery Grang (A3D3) under NSF grant #PHY-2117997.
Appendix A Methods
A.1 Data augmentations and training input
Hard-scattering events for $\text{p}\text{p}\to\text{j}\text{Z}$ and $\text{p}\text{p}\to\text{H}\text{Z}$, with $\text{H}\to\text{b}\bar{\text{b}}$ and $\text{Z}\to\nu\bar{\nu}$, from which we sample the jets for training the networks, are generated at a center-of-mass energy of $\sqrt{s}=13\,\text{TeV}$ using MadGraph5_aMC@NLO v3.4.0 [49] with the NNPDF30_nlo_nf_5_pdfas parton distribution function [50] at leading-order accuracy in QCD. The j and, respectively, the Higgs boson are generated within a $p_{\mathrm{T}}$ range of 400-600 GeV and within |$\absolutevalue{\eta}<0.1$. The hard-scattering events are then passed through a parton shower simulation and are reconstructed using the Delphes v3.4.3pre1 detector simulation [51] with a CMS detector-like geometry. For gluons and the Higgs, the first splitting or, respectively, decay is generated at matrix-element level with a minimum $\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}>0.001$. All further splittings are generated by the subsequent parton shower.
The RS3L space is trained with 5M events using with a 50%/50% admixture of QCD jets and Higgs jets. During the training of the network, the gradients get updated after each minibatch consisting of 200 jet samples. For each nominal jet, a systematic variation of the jet is included in the same minibatch, selected at random among the four possible variations at random to complete the (anchor, augmentation) pairs. The composition of nominal jets in the training set is always exactly 50%, while the remaining 50% is allocated to the augmentations used in training.
A.2 Network architecture
$h_{\mathrm{embed}}=\mathrm{MLP}_{\mathrm{embed}}(X)$
$h_{\mathrm{DGC1}}=\textsc{DynamicEdgeConv}(h_{\mathrm{embed}}|k=24)$
$h_{\mathrm{DGC2}}=\textsc{DynamicEdgeConv}(h_{\mathrm{DGC1}}|k=24)$
$h_{\mathrm{DGC3}}=\textsc{DynamicEdgeConv}(h_{\mathrm{DGC2}}|k=24)$
$h_{\mathrm{enc}}=\mathrm{MLP}_{\mathrm{enc}}(h_{\mathrm{DGC3}})$
$\bm{z}=\textsc{GlobalSumPool}(h_{\mathrm{enc}})\ .$ (1)
The input feature set $X$ is given by 14 features per particle, listed in Tab. 5. These input features are not normalized over a minibatch, but their magnitudes are comparable among the features and $\mathcal{O}(1)$, thereby regularizing the network and leading to a similar effect as a pre-transformation that enforces a mean of 0 and standard deviation of 1.
For embedding the 14 particle features into a higher-dimensional latent space, two fully-connected layers with dimensions (128/128) are used in $\textsc{MLP}_{\text{embed}}$. After a series of three convolutions with $k=24$, $\mathrm{MLP}_{\mathrm{enc}}$ consists of four fully-connected layers with dimensions (64/32/32/8). Outputs of all neurons in both MLPs are activated with an exponential linear unit [56]. The final building block, GlobalSumPool, aggregates the eight features of each particle in the jet from the last layer of $\mathrm{MLP}_{\mathrm{enc}}$ to obtain the final eight features $\bm{z}$ characterizing the jet in the RS3L space.
Table 5: Description of the particle input features.
Particle feature Description
$\log{p_{\text{T}}}$ ) logarithm of the transverse momentum
$\log{E}$ ) logarithm of the energy
$\log{(E/E^{\text{Jet}})}$ ) logarithm of $E$ normalized w.r.t. the jet
$\Delta\eta$ difference in pseudorapidity w.r.t. the jet
$\Delta\phi$ difference in $\phi$ w.r.t. the jet
$\Delta R$ distance in $\eta$-$\phi$ plane from the jet axis
$\tanh d_{0}$ hyperbolic tangent of the transverse impact parameter
$\tanh d_{z}$ hyperbolic tangent of the longitudinal impact parameter
isPhoton 1 if particle is a photon, else 0
isMuon 1 if particle is a muon, else 0
isElectron 1 if particle is an electron, else 0
isCH 1 if particle is a charged hadron, else 0
isNH 1 if particle is a neutral hadron, else 0
The network will try to push apart dissimilar jets (negative pairs) on the $N-1$-dimensional hypersphere with unit radius given by the $N$ input features, while trying to align similar jets (positive pairs) and characterize them by similar features. The competition between these two effects is regulated by a temperature parameter $\tau$. We studied this parameter based on the performance obtained when using the RS3L space to distinguish Higgs from QCD jets. Specifically, we trained RS3L spaces with $\tau$ ranging from $0.05$ to $0.5$ and fine-tuned these networks on the Higgs vs. QCD classification task with a fixed backbone. A fixed backbone offers the best insight into the performance of the RS3L pre-training step because it projects jets into a one-dimensional classification space using only the information derived during the self-supervised step. We observed that the Higgs vs. QCD binary classification performance was optimal with $\tau=0.1$. The full set of observed features is shown included in Figure 7.
Figure 7: The eight contrastive features derived during the self-supervision step.
The performance of the RS3L space fine-tuned on the in-domain Higgs vs. QCD classification task is shown in ROC curves Figure 9. A separate ROC curve is shown for Higgs separation from each individual QCD subprocesses. Note that the training sample consists of an admixture with similar weightings of all QCD subprocesses. The left (right) shows a graph with floating (fixed). Superior performance is observed for a graph with floating weights.
Figure 9: ROC curves showing the tagging performance of RS3L fine-tunings on the in-domain Higgs vs. QCD classification task. Left: graph fine-tuned with floating weights. Right: graph fine-tuned with fixed weights.
A.3 Fine-tuning and fully-supervised trainings
The MLP placed on top of the RS3L architecture is given by one additional fully-connected layer with 8 input dimensions and a single sigmoid activation function at the output. The backbone network architecture used in the fine-tuning is the exact same as the fully-supervised. The only difference between the approaches is that in the fine-tuning method the base graph trained with RS3L is loaded as a hot-start for the classification and the MLP is added on top of the backbone network. We consider two configurations of fine-tuning the RS3L space: a ``fixed'' graph (where only the MLP parameters are floating) and a ``floating'' graph (where all the MLP and graph parameters are floating). The admixture of QCD/H/W, as well as augmentations composition, are always consistent between the RS3L fine-tuning and full supervision. The RS3L space and classification steps are trained on independent datasets.

### v2

@@ III.2.1 In-distribution classification and robustness
Figure 6: Tagging performance on the Higgs vs. QCD classification task. All numbers are obtained on a sample comprising only jets from the default parton shower scenario. Fine-tunings on top of a RS3L space that was trained with 5M jets are shown in light and dark red, for graphs that have floating and fixed weights, respectively. The number in parentheses indicates the fine-tuning dataset size. Fully-supervised networks with various training sizes are shown in shades of blue. Three repeated trainings with different random initialization are used to compute an average and standard deviation for each ROC. These are shown as bands. Note that fixed-weight trainings do not have any appreciable uncertainty.
@@
The most straightforward application of RS3L is a binary classification fine-tuning to separate the same jet classes that are used to train the RS3L space. Thus, we first consider the aforementioned canonical classification task of discriminating between jets originating from Higgs bosons and jets originating from QCD.
@@
Two variants are considered for fine-tuning the RS3L space: (i) The weights of the backbone model are kept fixed, and a task-specific head consisting of a one-layer multilayer perceptron (MLP) is added, whose parameters are trained on the task at hand. (ii) The weights of the backbone model are used as initial parameters but are allowed to be adjusted (“floated”) together with the ones of the task head.
@@
A comparison of RS3L spaces fine-tuned for a Higgs vs. QCD classification task is presented in Figure 6, which shows the receiver operating characteristic (ROC) curve calculated from nominal jets only. The QCD rejection rate (1/ false positive rate) is shown in Table 2 for various Higgs tagging efficiencies, again computed from only nominal jets. Compared to a fully-supervised network trained on 8M jets, a fine-tuning using 3M jets on top of a 5M RS3L space yields a similar performance. If we instead match the yields of the training dataset of the fully-supervised approach and the fine-tuning (i.e., both training on 3M labeled samples), a gain in performance is observed for the RS3L + fine-tuning strategy. This indicates that a common, large pre-training, for which RS3L can be one possible strategy, allows for efficient fine-tuning on comparably much smaller dataset sizes than a fully supervised setup due to the “hot start”.
@@
Furthermore, it is important to note that all training configurations that use floating networks are more performant than networks trained with fixed weights on top of the RS3L backbone. This is to be expected because in the fixed weight case the model is severely restricted, as the only trainable parameters belong to the MLP head. Lastly, all networks outperform $N_{2}$ by a wide margin.
The performance of the RS3L space fine-tuned on the in-domain Higgs vs. QCD classification task is shown in ROC curves Figure 7 with separate ROC curves for each individual QCD subprocess. Note that the training sample consists of an admixture with similar weightings of all QCD subprocesses. The top (bottom) shows a graph with floating (fixed) weights. Superior performance is observed for a graph with floating weights in particular for light flavor QCD jets. This is an interesting observation, as it seems that the handles on flavor tagging such as the impact parameters of charged jet constituents get exploited during the fine-tuning step relative to the pre-training stage. For the fine-tuning with floating weights, all light-flavor subprocesses end up with roughly the same ROC curves, while for fine-tuning with fixed weights, the discrimination against single-quark light flavor jets (labelled “q”) is much better than against light-flavor gluon splittings (labelled “g(qq)”). In addition, the discrimination power against $\mathrm{g}\to\mathrm{b}\overline{\mathrm{b}}$ to first order does not stem from flavor information, as the Higgs boson also decays into a pair of b quarks. Hence, a direct comparison with the ROC curve for $N_{2}$ from Figure 6 allows an understanding of the substructure usage, with even fixed-weight RS3L improving over $N_{2}$ as much as 50% for a Higgs acceptance of 0.2.
Figure 7: ROC curves showing the tagging performance of RS3L fine-tunings on the in-domain Higgs vs. QCD classification task. Top: graph fine-tuned with floating weights. Bottom: graph fine-tuned with fixed weights.
@@
In addition to background rejection, it is necessary to examine the classifier under other metrics. In particular, we consider how the network response varies under different parton shower configurations used in the simulator (which we call the network’s “robustness”). The robustness is quantified by calculating the Wasserstein distance [53, 54, 55] between the one-dimensional distributions of the classifier output of nominal and, respectively, augmented jets as a measure of the similarity between the responses of the RS3L network to different inputs. The distances are presented in Table 3. Each column shows the distance between the distribution for augmented jets (indicated by column title) and nominal jets. A lower Wasserstein distance indicates a network that is less sensitive to the showering configurations. As expected, the distance between nominal jets and jets re-showered with just a different numerical seed is smaller than the distance between nominal jets and any other augmentation (“FSR up”, “FSR down”, or “Herwig”). Especially the distance between nominal jets and Herwig7-showered jets is significant.
@@
Comparing the different training strategies, the most robust networks are trained using the RS3L method. With a fixed-weight approach, the Wasserstein distance is significantly smaller than for all other networks. This is a direct benefit of the RS3L pre-training strategy of minimizing the distance between positive pairs. We note the significant trade-off in terms of worse QCD rejection rates (Table Table 2) for the fixed backbone. Including Herwig7 augmentations in RS3L or, respectively, full supervision greatly reduces the distances compared to setups without Herwig7 augmentations. Interestingly, we observe that already including seed and/or FSR augmentations in RS3L helps fine-tuned networks to be more robust against Herwig7-induced domain shift compared to fully supervised strategies. Whether this indicates that a gain in robustness is potentially attainable simply by training the backbone with re-seeded jets merits further investigation that we leave to future studies.
@@
Table 3: Wasserstein distances between distributions of tagger output for nominal and augmented jets. Each column indicates the distances between the nominal and the respective augmentation. The number of training events is included for each training setup. Fine-tunings are performed on a 5M RS3L pre-training with either fixed or floating backbone weights. The inclusion of specific augmentations (e.g. “seed” or “FSR”) in the training setup indicates that the pre-training and fine-tuning or supervision is performed on that set of augmentations. Uncertainties are derived from three independent trainings. The bolded numbers indicate the network with the smallest Wasserstein distances.
@@
III.2.2 Out-of-distribution classification task
We now shift to out-of-distribution tasks to address the degree to which a general representation of showering jets was achieved through the RS3L pre-training. To investigate this, we fine-tune RS3L for a classification task that falls outside of the distributions used for pre-training, namely the task of discriminating between QCD jets and jets from “hadronic” W boson decays. Contrary to the Higgs boson decays, jets from W boson decays do not feature $B$ hadrons [^2] The authors note that the absolute tagging performance on the W vs. QCD and Higgs vs. QCD datasets cannot be compared directly because of the composition of the sample, see section II.. The experimental signature therefore falls out of the distribution of the Higgs jets used to train the RS3L backbone.
@@ IV Outlook
Enforcing domain completeness through a strategy to embed the known modeling uncertainties within the self-supervised space directly maps this work to the quality of the simulation used at training. Improved simulators can lead to improved embeddings, which, in turn, lead to a better understanding of the underlying physics. As a result, RS3L can adapt to the next generation of simulation modeling that stands to be substantially better than the current simulation. Through self-supervised approaches, we can continually improve the quality of the physics that is extracted, provided we continue to improve the physics model in the samples used for the self-supervision.
In future work, alternative self-supervised learning approaches to the employed SimCLR strategy, such as VICReg [56], SimSiam [57], and BarlowTwins [58] should be explored to study the practical impact of precise components of their loss functions. Additionally, the dataset size for the pre-training should be varied to systematically study the regimes in which the performances of SSL, fine-tuning, and fully-supervised strategies may saturate. Finally, detailed comparisons should be made with other pre-training strategies proposed for HEP, such as masked particle modelling.
@@ V Acknowledgments
BM acknowledges the support of the Alexander von Humboldt foundation and of Schmidt Sciences. MK and JK are supported by the US Department of Energy (DOE) under grant DE-AC02-76SF00515. PH and JK are supported by the Institute for Artificial Intelligence and Fundamental Interactions (IAIFI) under the NSF grant #PHY-2019786 and the Accelerated AI Algorithms for Data Driven Discovery Grant (A3D3) under NSF grant #PHY-2117997.